# Less is More: Hop-Wise Graph Attention for Scalable and Generalizable Learning on Circuits

### 06 Mar 2024 v2

@@ III-C Complexity analysis of HOGA
Given a graph with $n$ nodes and $m$ edges, suppose the feature dimension is $d\ll n$ and there are $K\ll n$ hops considered, the complexity of hop-wise feature generation is $O(Kmd)$. Besides, the gated self-attention module has the complexity of $O(nKd^{2}+nK^{2}d)$ for performing linear projections and computing the self-attention matrix $S$ in Equation (7). Thus, the total complexity of HOGA is $O(Kmd+nKd^{2}+nK^{2}d)$, which is linear with respect to the number of nodes/edges.
IV Experiment
@@
Table 1: Statistics of OpenABC-D benchmark — The upper $20$ designs are used for training and the lower $9$ designs for evaluation.
@@
IV-A Experimental Setup
Table 2: Comparison of HOGA and GCN for QoR prediction — We choose mean absolute percentage error (MAPE) as the evaluation metric. Lower MAPE score means better QoR prediction. HOGA-$2$ and HOGA-$5$ indicate $K=2$ and $K=5$ in HOGA, respectively.
@@ IV-A Experimental Setup
Figure 4: GCN and HOGA-$5$ predictions vs. ground truth — The ground truth value is preprocessed in the OpenABC-D benchmark.
We evaluate HOGA on two representative EDA tasks: (1) QoR prediction. We focus on the OpenABC-D benchmark for predicting the optimized number of gates in And-Inverter-Graphs (AIGs) after logic synthesis optimization through ABC [3]. Notably, OpenABC-D consists of $870,000$ AIGs that are generated by running various synthesis recipes on $29$ open-source hardware IPs shown in Table 1, which contain different functionality designs from MIT LL labs CEP [5], OpenCores [1], and IWLS [2]. To demonstrate the generalizability of our approach, we train HOGA on the upper $20$ designs in Table 1 and evaluate it on the rest of the designs, which is a challenging graph-level regression task; (2) Functional reasoning. We follow Gamora [18] to identify adder blocks by predicting the sum and carry-out nodes in AIGs of carry-save array (CSA) and Booth multipliers, which is formulated as a node-level classification task. As is known that technology mapping can largely increase the reasoning complexity on Boolean networks [12, 19], we consider the most challenging scenario in Gamora, where AIGs are generated by ABC with complex ASAP 7nm technology mapping, and we only train HOGA on an $8$-bit multiplier design and perform inference on multipliers with bitwidth up to $768$.
Note that the authors of OpenABC-D and Gamora have proposed their own GNN-based methods for the aforementioned tasks respectively, which serve as our baselines. To ensure a fair comparison, we only replace their GNN models with HOGA and keep other components the same in both baseline methods, as depicted in Figure 3. In regard to hyperparameter settings of HOGA, we adopt the Adam optimizer with a learning rate of $0.0001$, a hidden dimension of $256$, and fix the number of gated self-attention layer to $1$. Besides, we set the number of hops $K$ as $5$ for experiments on OpenABC-D and $8$ on Gamora, which captures the information from the same number of hops as the baseline GNNs in both tasks. We leverage DistributedDataParallel in the PyTorch framework for distributed training on HOGA. All experiments are conducted on a Linux machine with an Intel Xeon Gold $5218$ CPU, which features 8 cores running at $2.30$ GHz, and is equipped with 4 RTX A6000600060006000 GPUs, each having 48 GB of memory.
IV-B Evaluation on QoR Prediction
Figure 5: Multi-GPU training time of HOGA on OpenABC-D.
Figure 6: Functional reasoning on AIG-based CSA and Booth multipliers with $7$nm technology mapping.
Our baseline is a $5$-layer GCN, as previously used for the OpenABC-D benchmark [6].
Figure 7: Visualization on hop-wise attention scores per node in the AIG-based $768$-bit Booth Multiplier — Rows and columns in each heatmap denote nodes and their hop-wise neighbors, respectively.
@@ IV-B Evaluation on QoR Prediction
Besides, we choose mean absolute percentage error (MAPE) as the evaluation metric, which is defined as $\text{MAPE}=\frac{1}{g}\sum_{i=1}^{g}|\frac{y_{i}-\hat{y}_{i}}{y_{i}}|\times 100\%$, where $y_{i}$ and $\hat{y}_{i}$ denote the ground truth and model prediction on the $i$-th sample (graph), respectively. Table 2 indicates that both HOGA models significantly outperform GCN across all test designs. In particular, HOGA-$5$ improves the estimation error over GCN on vga_lcd by a margin of $46.76\%$. The superiority of HOGA is further demonstrated in Figure 4, where the QoR predictions by HOGA-$5$ are highly correlated with the ground truth. In contrast, the GCN model fails to accurately predict the actual QoR values on those unseen designs. Moreover, by comparing HOGA-$5$ and HOGA-$2$, we can see the trade-off between the accuracy and training time of HOGA. Compared to the GCN baseline, HOGA-$2$ not only improves the average estimation error by $18.2\%$, but also achieves a notable 3.1×3.1$\times$3.1 × training speedup on a single GPU.
@@
It is noteworthy that HOGA can be easily accelerated through distributed training, owing to its high parallelism for learning node representations. As shown in Figure 5, the training time of HOGA almost linearly decreases when the number of GPU devices is increased. As a result, it takes only $3.5$ ($1.1$) hours to train HOGA-$5$ (HOGA-$2$) on the OpenABC-D benchmark. Notably, the first phase of HOGA for generating hop-wise features takes $13$ minutes, which is negligible compared to the overall training time. We believe that HOGA can be further accelerated if more computing resources are available, rendering it applicable to industrial-scale circuit applications.
IV-C Evaluation on Functional Reasoning
@@
To further demonstrate the generalization capability, we evaluate HOGA on the functional reasoning task [18] by identifying root nodes of XOR and majority (MAJ) operations in AIG-based multipliers, which correspond to sum and carry-out of adder blocks, respectively. More concretely, there are $4$ node categories for classification: (1) MAJ nodes; (2) XOR nodes; (3) nodes shared by MAJ and XOR; (4) other plain nodes such as PI/PO. Apart from choosing the GraphSAGE model used in Gamora as our baseline, we also consider GraphSAINT [20], a popular sampling-based GNN model, and SIGN [7] that adopts an MLP model on hop-wise features. Note that all baselines as well as HOGA are trained on an AIG-based $8$-bit multiplier and evaluated on multipliers with bitwidths ranging from $64$ to $768$.
@@
As shown in Figure 6, all baselines perform poorly on CSA multipliers, which indicates their limited generalizability to complex designs. In addition, GraphSAINT performs even worse than the vanilla GraphSAGE model on Booth multipliers, indicating sampling-based GNNs are not suitable for circuit graphs, as analyzed in Section II-A. While SIGN achieves reasonable accuracy on Booth multipliers, it still largely lags behind HOGA on CSA multipliers. We attribute it to our proposed gated self-attention module, which adaptively learns high-order features from different hops per node. This is further confirmed in Section IV-D. As a result, HOGA largely surpasses baseline models on both CSA and Booth multiplier circuits, with a $10.0\%$ accuracy improvement on the $768$-bit CSA multiplier. More importantly, Figure 6 illustrates that the accuracy of HOGA exhibits a rising trend as the bitwidth of test multipliers increases, which is crucial for accurate reasoning on large-scale Boolean networks.
IV-D Visualization on HOGA Attention Scores
@@
For the sake of clear visualization, we randomly sample $100$ nodes per classification category from the AIG of $768$-bit Booth multiplier, based on which four heatmaps are drawn in Figure 7. Notably, the rows and columns in each heatmap correspond to nodes and their hop-wise neighbors, respectively. Each element in a heatmap row represents the hop-wise attention score $c_{k}$ defined in Equation (10), which is used to adaptively combine $\hat{H}_{k}$ for producing the final node representations. Figure 7 clearly shows that HOGA is able to identify critical features from different hop $k$ for learning the underlying Boolean functions. Notably, since we use a single gated self-attention layer, each $\hat{H}_{k}$ captures second-order interactions among hop-wise features, corresponding to second-order graph structures. As a consequence, HOGA skips odd-hop neighbor features and primarily focuses on $\hat{H}_{k}$ with $k\in\{2,4,6\}$ for learning MAJ/XOR functions, as indicated by the attention scores shown in the first three heatmaps. As for the last heatmap, the attention scores of HOGA are relatively random since those nodes are PI/PO or plain AND gates, whose Boolean functions are not meaningful for this task.
V Conclusion
@@
This work introduces HOGA, a hop-wise attention approach for learning circuit representations in a scalable and generalizable way. HOGA first computes hop-wise features, which are then independently fed into a gated self-attention module for capturing high-order interactions among different hop neighbors. Experimental results showcase that HOGA not only comfortably scales to large-scale circuits via distributed training, but also outperforms conventional GNNs for generalizing to unseen and complex circuit designs.
Acknowledgments
@@
This work is supported in part by a Qualcomm Innovation Fellowship, NSF Awards #2118709 and #2212371, and ACE, one of the seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA.

### 10 Apr 2024 v4

@@ 3.3. Complexity analysis of HOGA
Suppose we consider $h$ feature dimensions and $K$ hops. The complexity of hop-wise feature generation is $O(Kh|\mathcal{E}|)$. Besides, the gated self-attention module has a complexity of $O(Kh^{2}|\mathcal{V}|+K^{2}h|\mathcal{V}|)$ for linear projections and computing Equation (5). Therefore, the total complexity of HOGA is $O(Kh|\mathcal{E}|+Kh^{2}|\mathcal{V}|+K^{2}h|\mathcal{V}|)=O(|\mathcal{E}|+|\mathcal{V}|)$, which is linear with respect to the number of nodes/edges.
4. Experiment
4.1. Experimental Setup
We evaluate HOGA on two tasks: (1) QoR prediction. We focus on the OpenABC-D benchmark for predicting the optimized gate count in And-Inverter-Graphs (AIGs) after logic synthesis optimization through ABC (brayton2010abc, ). Notably, OpenABC-D consists of $870,000$ AIGs that are generated by running various synthesis recipes on IPs from MIT LL labs CEP (brendon2019cep, ), OpenCores (opencore, ), and IWLS (albrecht2005iwls, ). To demonstrate the generalizability of our approach, we train HOGA on the top $20$ designs in Table 1 and evaluate it on the rest of the designs; (2) Functional reasoning. We follow Gamora (wu2023gamora, ) to identify adder blocks by predicting the sum and carry-out nodes in AIGs of carry-save array (CSA) and Radix-$4$ Booth multipliers. As technology mapping can largely increase functional reasoning complexity (li2013wordrev, ; yu2016formal, ), we consider the most challenging scenario in Gamora, where AIGs are generated by ABC with complex ASAP $7$nm technology mapping, and we only train HOGA on an $8$-bit multiplier design and perform inference on multipliers with bitwidth up to $768$.
Note that the authors of OpenABC-D and Gamora have proposed their own GNN-based models for the aforementioned tasks. To ensure a fair comparison, we only replace their GNN blocks with HOGA and keep other model components the same as shown in Figures 3(b) and 3(c). In regard to HOGA hyperparameter settings, we adopt Adam optimizer with a learning rate of $0.0001$, a hidden dimension of $256$, and fix the number of gated self-attention layer to $1$. Besides, we set the number of hops $K$ as $5$ for experiments on OpenABC-D and $8$ on Gamora, which captures the information from the same number of hops as the baseline GNNs in both tasks. We leverage DistributedDataParallel in PyTorch for distributed training on HOGA. All experiments are conducted on a Linux machine with an Intel Xeon Gold $5218$ CPU and 4 RTX A6000600060006000 GPUs.
@@ 4.1. Experimental Setup
Table 1. Statistics of OpenABC-D benchmark — The top and bottom designs are used for training and test, respectively.
@@
Table 2. Comparison of HOGA and GCN for QoR prediction — We choose mean absolute percentage error (MAPE) as the evaluation metric (Lower score is better). HOGA-$2$ and HOGA-$5$ indicate $K=2$ and $K=5$ in HOGA, respectively.
@@
Figure 4. GCN and HOGA-$5$ predictions vs. ground truth — The ground truth value is preprocessed in the OpenABC-D benchmark.
4.2. Evaluation on QoR Prediction
@@
Our baseline is a $5$-layer GCN, as previously used for the OpenABC-D benchmark (chowdhury2021openabc, ). Besides, we choose mean absolute percentage error (MAPE) as the evaluation metric, which is defined as $\text{MAPE}=\frac{1}{g}\sum_{i=1}^{g}|\frac{y_{i}-\hat{y}_{i}}{y_{i}}|\times 100\%$, where $y_{i}$ and $\hat{y}_{i}$ denote the ground truth and model prediction on the $i$-th sample (graph), respectively. Table 2 indicates that both HOGA models significantly outperform GCN across all test designs. In particular, HOGA-$5$ improves the estimation error over GCN on vga_lcd by a margin of $46.76\%$. The superiority of HOGA is further demonstrated in Figure 4, where the QoR predictions by HOGA-$5$ are highly correlated with the ground truth. In contrast, the GCN model fails to accurately predict the actual QoR values on those unseen designs. Moreover, by comparing HOGA-$5$ and HOGA-$2$, we can see the trade-off between the accuracy and training time of HOGA. Notably, HOGA-$2$ not only improves the average estimation error over GCN by $18.2\%$, but also achieves a notable 3.1×3.1$\times$3.1 × training speedup on a single GPU.
Figure 5. Multi-GPU training time of HOGA on OpenABC-D.
@@
It is noteworthy that HOGA can be easily accelerated through distributed training, owing to its high parallelism for learning node representations. As shown in Figure 5, the training time of HOGA almost linearly decreases when the number of GPU devices is increased. As a result, it takes only $3.5$ ($1.1$) hours to train HOGA-$5$ (HOGA-$2$) on the OpenABC-D benchmark. Notably, the first phase of HOGA for generating hop-wise features takes $13$ minutes, which is negligible compared to the overall training time. We believe that HOGA can be further accelerated if more computing resources are available, rendering it applicable to industrial-scale applications.
Figure 6. Functional reasoning on AIG-based CSA and Radix-$4$ Booth multipliers with $7$nm technology mapping.
4.3. Evaluation on Functional Reasoning
Figure 7. Visualization on hop-wise attention scores per node in the AIG-based $768$-bit Radix-$4$ Booth Multiplier — Rows and columns in each heatmap denote nodes and their hop-wise neighbors, respectively.
@@ 4.3. Evaluation on Functional Reasoning
To further demonstrate the generalization capability, we evaluate HOGA on the functional reasoning task (wu2023gamora, ) by identifying root nodes of XOR and majority (MAJ) operations in AIG-based multipliers, which correspond to sum and carry-out of adder blocks, respectively. More concretely, there are $4$ node categories for classification: (1) MAJ nodes; (2) XOR nodes; (3) nodes shared by MAJ and XOR; (4) other plain nodes such as PI/PO. Apart from choosing GraphSAGE in Gamora as our baseline, we also consider GraphSAINT (zeng2019graphsaint, ), a popular sampling-based GNN model, and SIGN (frasca2020sign, ) that adopts an MLP model on hop-wise features. Note that all baselines as well as HOGA are trained on an AIG-based $8$-bit multiplier and evaluated on multipliers with bitwidths ranging from $64$ to $768$.
@@
As shown in Figure 6, all baselines perform poorly on CSA multipliers, which indicates their limited generalizability to complex designs. In addition, GraphSAINT performs even worse than GraphSAGE on Booth multipliers, indicating sampling-based GNNs are not suitable for circuit graphs, as analyzed in Section 2.1. While SIGN achieves reasonable accuracy on Booth multipliers, it still largely lags behind HOGA on CSA multipliers. We attribute it to our proposed gated self-attention module, which adaptively learns high-order features from different hops per node. This is further confirmed in Section 4.4. As a result, HOGA largely surpasses baseline models on both CSA and Booth multiplier circuits, with a $10.0\%$ accuracy improvement on the $768$-bit CSA multiplier. More importantly, Figure 6 illustrates that the accuracy of HOGA exhibits a rising trend as the bitwidth of test multipliers increases, which is crucial for accurate reasoning on large-scale Boolean networks.
4.4. Visualization on HOGA Attention Scores
@@
For the sake of clear visualization, we randomly sample $100$ nodes per classification category from the AIG of $768$-bit Booth multiplier, based on which four heatmaps are drawn in Figure 7. Notably, the rows and columns in each heatmap correspond to nodes and their hop-wise neighbors, respectively. Each element in a heatmap row represents the hop-wise attention score $c_{k}$ defined in Equation (6), which is used to adaptively combine important hop-wise features. Figure 7 clearly shows that HOGA is able to identify critical features from different hop $k$ for learning the underlying Boolean functions. Notably, since we use a single gated self-attention layer, each $\hat{H}_{k}$ in Equation (6) captures second-order interactions among hop-wise features, corresponding to second-order graph structures. As a consequence, HOGA skips odd-hop neighbor features and primarily focuses on $\hat{H}_{k}$ with $k\in\{2,4,6\}$ for learning MAJ/XOR functions, as indicated by the attention scores shown in the first three heatmaps. As for the last heatmap, the attention scores of HOGA are relatively random since those nodes are PI/PO or plain AND gates, whose Boolean functions are not meaningful for this task.
5. Conclusion
@@
This work introduces HOGA, a hop-wise attention approach for scalable and generalizable circuit representation learning. HOGA precomputes hop-wise features that are then fed into a gated attention module for capturing critical circuit structures. Our results showcase that HOGA not only comfortably scales to large-scale circuits via distributed training, but also outperforms conventional GNNs for generalizing to unseen and complex circuit designs.
Acknowledgements.
@@
This work is supported in part by a Qualcomm Innovation Fellowship, NSF Awards #2047176, #2118709 and #2212371, and ACE, one of the seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA.